If you wish to produce 45.0 nm x-rays in the laboratory, what is the minimum voltage you must use in accelerating the electrons?


1
Expert's answer
2020-06-11T10:44:54-0400

Law of conservation of energy is what can help us to solve the problem:


"qV=\frac{hc}{\lambda},\\\space\\
V=\frac{hc}{q\lambda}=\frac{(6.6\cdot10^{-34})(3\cdot10^8)}{(1.6\cdot10^{-19})(450\cdot10^{-9})}=2.8\text{ V}."